OpenStudy (anonymous):

HELP!! Solve the inequality 3x + 4y > 12 for y. Do not graph. If you just give me the anwser I'll give u a medel

OpenStudy (tyteen4a03):

@kiamonstaa 's answer is not correct. This is an inequality, not an equation. To make y the subject of the inequality, subtract 3x from both sides, then divide 4 on both sides.

OpenStudy (anonymous):

3x+4y>12 => 3x +4y = 12, so y = (-3x +12)/4 So y > ( -3x +12 )/4

OpenStudy (tyteen4a03):

@Frank1991 An inequality CANNOT be turned into an equation.

OpenStudy (anonymous):

yes it can, just turn it back in your last step. You could check it with an random number for x

OpenStudy (anonymous):

y>3-0.75x

OpenStudy (tyteen4a03):

@Frank1991 No. Although the way of solving inequalities are similar to solving equations, you cannot say that an inequality is an equation, because while equation have finite answers, inequalities have an infinite answer set.

OpenStudy (anonymous):

@tyteen4a03 i understand your concern, but i think it is easier to turn it to an equality because it is familiar. And just check in your last step in the unequal sign is faced the correct way.

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

i think my answer is correct.

OpenStudy (tyteen4a03):

@Frank1991 Mathematics does not allow you to turn something into another thing just because it's "more familiar". You have to give proof. While it's OK for you to relate those two, you cannot say that an inequality is an equation, that is not an acceptable behaviour in mathematics.

OpenStudy (anonymous):

The answer is y > -3/4x +3. and yes @kiamonstaa is correct

OpenStudy (anonymous):

@kiamonstaa can you please post hoe you got that?

OpenStudy (tyteen4a03):

@hali12 I just explained... To make y the subject of the inequality, subtract 3x from both sides, then divide 4 on both sides. 3x + 4y > 12 4y > 12 - 3x y > (12-3x)/4 y = 3 - 3/4x

OpenStudy (anonymous):

well...yea ha

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

welcome